The quality manager of a factory wants to determine the life expectancy for manufactured lamps of his factory and found the mean is 375 hours. The standard deviation of the population is 100 hours. A sample of 64 lamps showed a mean life time of 350 hours. At the 5% level, prove that the mean life expectancy is less than 375 hours. 


1
Expert's answer
2022-01-17T18:52:32-0500

The following null and alternative hypotheses need to be tested:

"H_0:\mu\geq375"

"H_1:\mu<375"

This corresponds to a left-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05,"

and the critical value for a left-tailed test is "z_c = -1.6449."

The rejection region for this left-tailed test is "R = \{z: z < -1.6449\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{350-375}{100/\sqrt{64}}=-2"

Since it is observed that "z = -2 <-1.6449= z_c ," it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value is "p=P(Z<-2)=0.02275,"  p and since "p = 0.02275 < 0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is less than "375," at the "\alpha = 0.05" significance level.

Therefore, there is enough evidence to claim that the mean life expectancy is less than 375 hours, at the "\alpha = 0.05" significance level.